# Progressive Quantum Algorithm for Maximum Independent Set with Quantum Alternating Operator Ansatz

### v2

@@ Abstract
Recently, Hadfield et al. proposed the Quantum Alternating Operator Ansatz (QAOA+) to tackle Constrained Combinatorial Optimization Problems (CCOPs). This paper proposes a Progressive Quantum Algorithm (PQA) to reduce the required qubits in solving the Maximum Independent Set (MIS) problem using QAOA+ ansatz. PQA constructs a subgraph that contains the MIS solution of the target graph $G$ and then solves the MIS problem on this subgraph to obtain the solution for $G$. To induce such a subgraph, PQA starts with a small-scale initial subgraph and progressively expands its graph size utilizing designed expansion rules. After each expansion, PQA solves the MIS problem on the current subgraph using QAOA+. In each run, PQA repeats the graph expansion and solving process until a predefined stopping condition is reached. Simulation results demonstrate that to achieve an optimal approximation ratio of 0.95, PQA requires only $5.5565\%$ ($15.4\%$) of the qubits and $10.695\%$ ($7.23\%$) of the runtime compared with QAOA+ on Erdős-Rényi (regular) graphs, highlighting the efficiency of PQA.
@@ I INTRODUCTION
Over recent years, quantum computing has emerged as a promising computational paradigm, demonstrating its potential for addressing complex problems such as database searching Grover (1997), prime factorization Shor (1999), data dimension reduction Lloyd et al. (2014); Yu et al. (2019); Pan et al. (2020); Wan et al. (2018) and solving equations Harrow et al. (2009). However, the practical implementation of these quantum algorithms is hindered by noise and the limited qubit capacity of existing Noisy Intermediate-Scale Quantum (NISQ) devices Preskill (2018); Cerezo et al. (2021); Bharti et al. (2022). Therefore, hybrid quantum-classical algorithms Grimsley et al. (2019); Tang et al. (2021); Zhu et al. (2019); Du et al. (2022); He et al. (2024); Song et al. (2024); Wu et al. (2023a, b); Wang et al. (2021); Liu et al. (2021) have been proposed to harness the capabilities of NISQ devices more effectively. The Quantum Approximate Optimization Algorithm (QAOA) Farhi et al. (2014); Bravyi et al. (2020); Zhou et al. (2020); Sack and Serbyn (2021); Streif and Leib (2020); Zhu et al. (2022); Ni et al. (2024); Zhou et al. (2023); Wurtz and Lykov (2021) is a hybrid quantum-classical algorithm with broad applications. Specifically, it has been used to solve the exact cover Vikstål et al. (2020), vertex cover Zhang et al. (2022), and other Combinatorial Optimization Problems (COPs) Wurtz and Love (2021); Ruan et al. (2023); Brandhofer et al. (2022); Tomesh et al. (2023); Finžgar et al. (2023). Constrained COPs (CCOPs) are a crucial subclass of combinatorial optimization, and solving CCOPs requires not only finding the optimal solution but also satisfying specific constraints, making them more complex and challenging compared with Unconstrained COPs (UCOPs) Karp (2010). To address these challenges for solving CCOPs, Hadfield et al. Hadfield et al. (2019) proposed an improved version of the original QAOA Farhi et al. (2014), known as QAOA+. QAOA+ constructs a Hilbert space that only contains feasible solutions to the problem, thereby reducing the solution search space compared with the original QAOA Ruan et al. (2023). Currently, QAOA+ has been applied to solve Maximum Independent Set (MIS) Saleem (2020), minimum exact cover Wang et al. (2023) and other CCOPs Hadfield et al. (2017); Wang et al. (2020); He et al. (2023). Given the limited resources of NISQ devices, developing heuristic strategies to reduce the qubit requirements of QAOA+ in solving CCOPs becomes essential. Recently, Tomesh et al. Tomesh et al. (2022a) proposed Quantum Local Search (QLS) with the QAOA+ ansatz and studied its performance on the MIS problem. Different from directly solving the MIS problem using QAOA+ (DS-QAOA+) Saleem (2020), QLS constructs a global solution by iterating through many local subproblems. Specifically, QLS constructs multiple local subgraphs (that may be of indeterminate graph size) based on certain strategies and solves the MIS problem on these subgraphs using QAOA+ to obtain local solutions. These local solutions are combined to form a global solution using classical methods. For DS-QAOA+, several qubits equal to the size of the graph are required. By breaking the problem into smaller subgraphs, QLS reduces the number of qubits needed for the MIS problem. Inspired by the Proactively Incremental Learning (PIL) algorithm Li et al. (2023), a new method called the Progressive Quantum Algorithm (PQA) is proposed to reduce the required qubits for solving the MIS problem. PQA aims to construct a subgraph containing the MIS solution of the original problem and then solve the MIS problem on this subgraph to obtain the final solution. Specifically, PQA starts with a small, sparse initial subgraph and progressively expands its size following rules distinct from QLS. After each expansion, PQA solves the MIS problem on the newly generated subgraph using QAOA+. In each run, PQA repeats the graph expansion and solving process until a predefined stopping condition is reached. By doing so, PQA may solve the original MIS problem on a smaller-scale subgraph compared with DS-QAOA+, and it does not require reconstructing the global solution from multiple local solutions compared with QLS. In our simulations, we compare the total qubits and runtime consumed by PQA, QLS, and DS-QAOA+ on multiple Erdős-Rényi (ER) and $k(=2,3)$-regular graphs when they achieve the same optimal approximation ratio (OAR). Across all tested graphs, PQA usually achieves the same OAR with fewer qubits and runtime than DS-QAOA+ and QLS. Specifically, for an OAR of 0.95, PQA uses $5.5565\%$ ($26.77\%$) of the qubits required by DS-QAOA+ (QLS) on ER graphs, and $15.4\%$ ($52.605\%$) on regular graphs, respectively. Additionally, PQA reaches the same OAR in $10.695\%$ ($75.365\%$) of the runtime required by DS-QAOA+ (QLS) on ER graphs, and $7.23\%$ ($63.91\%$) on regular graphs. The MIS problem is one of the most widely studied CCOPs, partly due to its broad applicability in various domains and its equivalence to the minimum vertex cover and maximum clique. The efficiency of PQA in solving the MIS problem not only may enhance the practicality of quantum algorithms but also pave the way for solving larger-scale problems in the future. This paper is structured as follows: Section II serves as the foundational segment, providing background knowledge on the MIS problem, QAOA, and its extension, QAOA+. Following this, in Section III, we delineate the progressive quantum algorithm, emphasizing the construction rules of the desired derived subgraph. Subsequently, in Section IV, we present a comprehensive set of numerical simulation experiments aimed at showcasing the quantum resource efficiency of PQA over DS-QAOA+ and QLS. Finally, in Section V, we offer a succinct conclusion and discussion to encapsulate the key findings of our study and the next research direction.
@@ II PRELIMINARIES
In this section, we introduce some preliminaries to help readers understand our work better.
@@ II.1 Review of MIS
MIS is defined on a graph $G=(V,E)$, where $V=\left\{{0,...,n-1}\right\}$ represents the set of vertices, and $E=\left\{\left(u,v\right)\right\}$ represents the set of edges, with the two vertices $u$ and $v$ on the edge $(u,v)$ being adjacent. In this paper, we denote the number of vertices and edges as $n$ and $m$, respectively. A vertex subset $V_{s}$ is called an independent set of $G$ if no two vertices in $V_{s}$ are adjacent. The MIS is the vertex subset with the largest number of vertices among all independent sets, and the number of vertices in the MIS is known as the independence number, denoted as $\beta(G)$ Karp (2010). While the solution to MIS for a graph $G$ may not be unique, the independence number is always determined. For any target graph $G$, the independent vertex subsets in an induced subgraph $G_{q}$ of $G$ are feasible solutions for $G$. There is at least one induced subgraph $G_{q}=(V_{q},E_{q})$ that has an identical MIS solution to $G$, where $V_{q}\subseteq V$ and $E_{q}\subseteq E$. As illustrated in FIG. 1, the right graph is the induced subgraph of the left graph, and they have the same MIS solution $\left\{1,2,3\right\}$, but the MIS solution $\left\{1,2,4\right\}$ to FIG 1 is not the MIS solution of FIG. 1 because the node $4$ is not in this induced subgraph. This implies that the MIS solution of the target graph may not always be a solution for its induced subgraph, and an independent set in $G$ may not be feasible for $G_{q}$.
@@
In this paper, we utilize $x_{u}=1$ to represent that the vertex $u$ is in the vertex subset, otherwise, $x_{u}=0$, where $u=0,1,...,n-1$. Each way of vertex partition corresponds to a unique bit string $x$, and there are $2^{n}$ forms for $x$ in total, where $x=x_{0}x_{1}\dots x_{n-1}$. The goal for MIS is to maximize the objective function
@@
when the independence constraint $\sum_{(u,v)\in E}x_{u}x_{v}=0$ is met. This constraint can be incorporated into the objective function using the Lagrange multiplier method Brady and Hadfield (2023), thus converting the constrained problem to unconstrained problem. The obtained classical cost function is
@@
where the value of $\lambda$ is determined through additional hyper-parameter optimization steps Finžgar et al. (2023), and $\lambda>1$ can impose penalties on those terms that violate independence constraints Lucas (2014), making the maximal cost function value correspond to the exact MIS solution instead of other feasible or infeasible bit strings. Here, $C_{\text{max}}$ equals the independence number $\beta(G)$.
@@ II.2 Review of QAOA
Inspired by quantum adiabatic evolution Albash and Lidar (2018), QAOA encodes the solution of the problem into the ground state of the target Hamiltonian $H_{C}$ and aims to gradually evolve towards this ground state through a parameterized quantum circuit (PQC). This evolution starts from the ground state $|s\rangle$ of the initial Hamiltonian $H_{B}$, with a quantum circuit built using $p$-level QAOA ansatz. Notably, the initial state $|s\rangle$ is required to be trivial to implement. That is, $|s\rangle$ can be prepared by a constant-depth quantum circuit. For the original QAOA, one level QAOA ansatz consists of two variational QAOA parameters (denoted as $\gamma_{i}$ and $\beta_{i}$) and two unitaries $e^{-i\gamma_{i}H_{C}}$ and $e^{-i\beta_{i}H_{B}}$, where $i=1,2,\cdots,p$, and $p$ is the level depth. Generally, the initial Hamiltonian is chosen as $H_{B}=-\sum_{u=0}^{n-1}\sigma_{u}^{x}$ because its ground state $|s\rangle=|+\rangle^{\otimes n}$ can be trivial to prepare. Here, $\sigma_{u}^{x}$ refers to the Pauli-X operator applied to the $u-$th qubit, effectively flipping its state. To encode the MIS solution into the ground state of $H_{C}$, the cost function $-C(x)$ is transformed into the target Hamiltonian
@@
by transforming each binary variable $x_{u}$ to a quantum spin $\frac{I-\sigma_{u}^{z}}{2}$ Farhi et al. (2014), where $\sigma_{u}^{z}$ refers that applying Pauli-Z to the $u$-th qubit. The ground state of $H_{C}$ can be approximately obtained by alternately applying unitaries operators $e^{-i\gamma_{i}H_{C}}$ and $e^{-i\beta_{i}H_{B}}$ to the initial quantum state $|s\rangle$. After applying these unitaries for $p$ levels (corresponding to $i=1,2,\cdots,p$), the output state of the PQC is
@@
where $\bm{\gamma_{p}}=(\gamma_{1},...,\gamma_{p})$ and $\bm{\beta_{p}}=(\beta_{1},...,\beta_{p})$ are $2p$ variational QAOA parameters. The expectation value $F(\bm{\gamma_{p}},\bm{\beta_{p}})$ of $H_{C}$ in this variational quantum state is defined as
@@
which can be calculated by repeated measurements of the output quantum state using the computational basis, and the goal is to search for the global optimal parameters
$(\bm{\gamma_{p}}^{opt},\bm{\beta_{p}}^{opt})=\arg\underset{(\bm{\gamma_{p}},\bm{\beta_{p}})}{\max}F(\bm{\gamma_{p}},\bm{\beta_{p}}).$ (6)
@@
The search for optimal parameters is typically done by starting with some guesses of the initial parameters, where the quantum computer computes the expectation value of the output state and delivers the relevant information to the classical computer. These parameters are optimized by the classical optimizer and then delivered to the quantum computer. The outer parameter optimization is repeated until it meets the termination condition (e.g., the maximal number of iterations or convergence tolerance $\varepsilon$ Ni et al. (2024)). To quantify the quality of the QAOA solution (i.e., measure the difference between the output state of QAOA and the ground state of $H_{C}$), we introduce the approximation ratio (AR)
$r=\frac{F({\bm{\gamma_{p}}},{\bm{\beta_{p}}})}{F_{\text{max}}},$ (7)
@@
where $F_{\text{max}}$ is the negative of the ground state energy of $H_{C}$, with equaling $C_{\text{max}}$. For small graph instances, $F_{\text{max}}$ can be obtained by brute force search. For large graphs, an approximation of $F_{\text{max}}$ can be obtained by executing multiple runs of the greedy algorithm Halldórsson and Radhakrishnan (1994) and selecting the largest independent set found across these runs. AR reflects how close the solution provided by QAOA is to the exact solution, $r\leq 1$, with the value of 1 as the exact solution.
II.3 Review of QAOA+
The original QAOA searches for the ground state of $H_{C}$ from a Hilbert space that includes both feasible and infeasible quantum states. As a result, the output states may not always satisfy the independence constraint and violate one or more constraints Ruan et al. (2023); Finžgar et al. (2023). To overcome this challenge, Hadfield et al. Hadfield et al. (2017, 2019) introduce a novel quantum alternating operator denoted as QAOA+. Different from incorporating the independence constraint into the target Hamiltonian utilizing the Lagrange multiplier, QAOA+ encodes constraints into the mixer Hamiltonian $H_{B}$. This allows QAOA+ to search for the optimal solution within a feasible subspace that only contains feasible quantum states. Consequently, the target Hamiltonian
$H_{C}=\sum_{u=0}^{n-1}\frac{\sigma_{u}^{z}-I}{2}$ (8)
is obtained by converting the $x_{u}$ in $-C_{1}(x)$ to a quantum spin $\frac{I-\sigma_{u}^{z}}{2}$, thereby encoding the MIS solution into the ground state of $H_{C}$. Similar to QAOA, QAOA+ approaches the ground state of $H_{C}$ by alternately applying unitaries $e^{-i\gamma_{i}H_{C}}$ and $e^{-i\beta_{i}H_{B}}$ on an initial quantum state $|s\rangle$ required to be trivial to implement. However, in QAOA+, the initial quantum state must be the superposition of feasible states. This requirement ensures that the feasible subspace contains no states that violate the constraints. In addition, the initial quantum state $|s\rangle$ may not be the ground state of $H_{B}$ in QAOA+, unlike in QAOA.
Figure 2: The implementation of $e^{-i\beta_{i}B_{u}}$ when $d(u)\geq 1$. The multi-qubit controlled RX gate can be decomposed into two multi-qubit controlled gates and a single-qubit controlled RX gate, utilizing an auxiliary qubit Vale et al. (2023). The auxiliary qubit is initialized as $|0\rangle$ and is flipped to $|1\rangle$ only when the quantum states of all adjacent vertices are $|0\rangle$ (i.e., $f(u)=0$). Once this condition is met, an RX gate is applied on the $u$-th qubit. After the operation, the state of the auxiliary qubit can be reset from $|1\rangle$ to $|0\rangle$ using the right multi-qubit controlled gate operation. Otherwise, the auxiliary qubit is always $|0\rangle$ when $f(u)>0$, and there is an Identity operation applied to the $u$-th qubit.
Figure 3: A schematic of directly solving the MIS problem on the target graph using QAOA or QAOA+ ansatz, where the concrete form of the initial quantum state $|s\rangle$, the mixer Hamiltonian $H_{B}$, and the target Hamiltonian $H_{C}$ are relevant to the choose of ansatz.
The design of the mixing Hamiltonian $H_{B}$ must preserve the feasibility of the subspace (i.e., ensure the subspace only has quantum states obeying constraints) and explore a new feasible subspace (i.e., provide transitions between various feasible states) Hadfield et al. (2019). For an independent vertex subset $V_{s}$, deleting an arbitrary vertice $v_{a}$ (i.e., the state of $x_{v_{a}}$ is converted from 1 to 0) from $V_{s}$ does not change its independence. However, a new vertex $v_{a}$ can be added to $V_{s}$ (i.e., flipping $x_{v_{a}}$ from 0 to 1) without destroying its independence only when there are no adjacent vertices of $v_{a}$ in $V_{s}$. Thus, for any node, controlling its bit-flip operation based on the states of its neighbors is sufficient to both add and remove vertices while preserving the independence property Hadfield et al. (2019). The bit-flip of $v_{a}$ occurs only if
$f(v_{a})=\sum_{j=1}^{d(v_{a})}x_{v_{j,a}}=0,$ (9)
where $v_{j,a}$ is the $j$-th adjacent vertex of $v_{a}$, and $d(v_{a})$ is the number of adjacent vertices of $v_{a}$, with $j=1,\cdots,d(v_{a})$. Mapping to a quantum circuit, for any vertex $u$, a Pauli-X operation is applied on the $u$-th qubit (i.e., this qubit is flipped) if and only if all quantum states of adjacent vertices of $u$ are $|0\rangle$ (i.e., $f(u)=0$). Otherwise, a Pauli-I operation is applied on the $u$-th qubit. The corresponding partial mixer Hamiltonian $B_{u}$ can be described as
The mixer Hamiltonian is given by $H_{B}=\sum_{u=0}^{n-1}B_{u}$. Specifically, the quantum state resulting from applying $e^{-i\beta_{i}B_{u}}$ to the single feasible quantum state $|x_{o}\rangle$ is
$e^{-i\beta_{i}B_{u}}|x_{o}\rangle=\begin{cases}&RX(2\beta_{i})|x_{o}\rangle,f(u)=0\\ &|x_{o}\rangle,f(u)>0,\end{cases}$ (11)
where the detailed result of $RX(2\beta_{i})|x_{o}\rangle$ is $\cos 2\beta_{i}|x_{o}\rangle-i\sin 2\beta_{i}|x_{0}\cdots\bar{x_{u}}\cdots x_{n-1}\rangle$, and $\bar{x_{u}}$ denotes the state of $u$-th qubit is converted from $|0\rangle$ to $|1\rangle$ (or ($|1\rangle$ to $|0\rangle$), implementing the addition or removal of node $u$ without generating infeasible quantum states. The quantum circuit for $e^{-i\beta_{i}B_{u}}$, shown in FIG. 2, corresponds to a multi-qubit controlled RX gate when the node has $d(u)\geq 1$ adjacent vertices, and to a single RX gate when $d(u)=0$. To help readers understand our subsequent work better, FIG. 3 provides a schematic illustrating the process of directly solving the MIS problem on the target graph using QAOA or QAOA+. Additionally, Table 1 is provided to clearly illustrate the differences between using QAOA and QAOA+ to solve the MIS problem.

### v3

@@ Abstract
Hadfield et al. proposed a novel Quantum Alternating Operator Ansatz algorithm (QAOA+), and this algorithm has wide applications in solving constrained combinatorial optimization problems (CCOPs) because of the advantages of QAOA+ ansatz in constructing a feasible solution space. In this paper, we propose a Progressive Quantum Algorithm (PQA) with QAOA+ ansatz to solve the Maximum Independent Set (MIS) problem using fewer qubits. The core idea of PQA is to construct a subgraph that is likely to contain the MIS solution of the target graph and then solve the MIS problem on this subgraph to obtain an approximate solution. To construct such a subgraph, PQA starts with a small-scale initial subgraph and progressively expands its graph size utilizing heuristic expansion strategies. After each expansion, PQA solves the MIS problem on the newly generated subgraph. In each run, PQA repeats the expansion and solving process until a predefined stopping condition is reached. Simulation results demonstrate that to achieve an approximation ratio of 0.95, PQA requires only $5.565\%$ ($2.170\%$) of the qubits and $17.59\%$ ($6.430\%$) of the runtime compared with directly solving the original problem using QAOA+ on Erdős-Rényi (3-regular) graphs, highlighting the efficiency of PQA.
@@ I Introduction
Quantum computing has made significant strides in recent years, showing great potential in solving complex problems such as database searching grover , prime factorization shor , data dimension reduction qpca ; reduction_yu ; reduction_pan ; reduction_wan , cross-validation Lijing , and solving equations hhl . Despite these promising advancements, the practical implementation of quantum algorithms faces challenges due to the noise and limited qubit capacity of current Noisy Intermediate-Scale Quantum (NISQ) devices NISQ ; VQA_review1 ; VQA_review2 . To fully harness the power of NISQ devices, hybrid quantum-classical algorithms adaptive-vqe ; qubit-adaptive-vqe ; QCL ; QAS ; RE-ADAPT-VQA ; QFL ; wu_quantum ; wu_qst have been proposed. The Quantum Approximate Optimization Algorithm (QAOA) qaoa ; RQAOA ; tree-QAOA ; adaptive-qaoa is a class of hybrid quantum-classical algorithms widely used to solve a variety of Combinatorial Optimization Problems (COPs) xiumei ; exact_cover ; INTERP ; perform1 ; constraints_QAOA ; portfolio ; divide_MIS ; MLI , demonstrating its versatility in addressing both unconstrained and constrained optimization tasks. In the context of Constrained COPs (CCOPs), QAOA typically encodes problem constraints into the target Hamiltonian by introducing penalty terms minimum_vertex_cover ; QIRO ; Ising_formulations and subsequently searches for the quasi-optimal solution using a Parameterized Quantum Circuit (PQC). However, this approach can result in infeasible solutions (i.e., solutions that violate one or more problem constraints) since QAOA explores the entire Hilbert space, including both feasible and infeasible states, without explicit constraints on the solution space RUAN . To overcome these challenges, Hadfield et al. introduced a novel Quantum Alternating Operator Ansatz algorithm (QAOA+) alternating_QAOA_ansatz , an enhanced version of the original QAOA. QAOA+ constructs a Hilbert space that includes only feasible solutions by encoding the problem constraints into the mixer Hamiltonian, and then searches for a quasi-optimal solution within this constrained space. This approach reduces the search space and eliminates the risk of infeasible solutions, offering a significant advantage over QAOA constraints_QAOA . QAOA+ has been successfully applied to many CCOPs graph_coloring1 ; graph_coloring2 ; portfolio_optimization ; MIS ; minimum_exact_cover , such as the Minimum Exact Cover problem. Although QAOA+ offers significant advantages in solving CCOPs, reducing the qubit requirements of QAOA+ is essential to tackle larger-scale CCOPs given the limited resources of current NISQ devices, thereby further expanding its applications. To achieve this goal, Tomesh et al. proposed a Quantum Local Search (QLS) algorithm with the QAOA+ ansatz and evaluated its performance on the Maximum Independent Set (MIS) problem QLS that is a significant CCOP encountered in various fields, including network design MIS_in_network and scheduling MIS_in_scheduling , and is equivalent to the minimum vertex cover or finding a maximum clique in the complementary graph MIS_vertex_cover . Unlike directly solving the MIS problem on the original graph using QAOA+ (DS-QAOA+) MIS , QLS constructs a global solution by solving many local subproblems. Specifically, QLS constructs multiple local subgraphs using specific strategies and solves the MIS problem on each subgraph with QAOA+ to obtain local solutions. These local solutions are then combined into a global solution through classical methods. By breaking the original problem into multiple sub-problems, QLS reduces the number of qubits needed for solving the MIS problem on the original graph, making it more suitable for current quantum devices with limited resources. Although the qubit requirements decrease as the subgraph size reduces, the complexity of merging local solutions increases as the number of subgraphs grows. In this paper, we propose a quantum heuristic algorithm called the Progressive Quantum Algorithm (PQA), inspired by the Proactively Incremental Learning algorithm PIL and the properties of the MIS problem, to reduce the qubit requirements of QAOA+ in solving the MIS problem. The core idea of PQA is to search for a subgraph that is likely to contain the MIS solution of the original graph and then solve the MIS problem on this subgraph to obtain an approximate solution. To construct such a subgraph, PQA commences with a small and sparse subgraph, gradually enlarging its size through strategies distinct from those utilized in QLS. Following each expansion, PQA employs QAOA+ to address the MIS problem on the freshly constructed subgraph. This process of graph expansion and problem resolution is iterated by PQA until a predetermined termination criterion is satisfied. By doing so, PQA can solve the original MIS problem in a smaller subgraph than DS-QAOA+. Additionally, unlike QLS, multiple local solutions are not required to be combined to reconstruct the global solution. Numerical results demonstrate that PQA generally achieves the same approximation ratio (AR) with fewer qubits and shorter runtime compared with DS-QAOA+ and QLS. Specifically, for an AR of 0.95, PQA uses $5.57\%$ of the qubits required by DS-QAOA+ and $26.77\%$ of those required by QLS on ER graphs, and $2.17\%$ and $21.76\%$ on 3-regular graphs, respectively. Additionally, PQA achieves the same AR in $17.59\%$ of the runtime required by DS-QAOA+ and $76.72\%$ of that required by QLS on ER graphs, and $6.43\%$ and $61.49\%$ on 3-regular graphs. The efficiency of PQA in solving the MIS problem can enhance the practicality of quantum algorithms and open up possibilities for solving large-scale problems in the future. This paper is structured as follows. Section II serves as the foundational segment, providing background knowledge on the MIS problem, QAOA, and its extension, QAOA+. Following this, in Section III, we delineate the progressive quantum algorithm, emphasizing the construction strategies of the desired derived subgraph. Subsequently, in Section IV, we present a comprehensive set of numerical simulation experiments aimed at showcasing the quantum resource efficiency of PQA over DS-QAOA+ and QLS. Finally, in Section V, we offer a succinct conclusion and discussion to encapsulate the key findings of our study and the next research direction.
@@ II Preliminaries
To help readers understand our work better, we first introduce the MIS problem in Section II.1, followed by an introduction to QAOA in Section II.2. Finally, we review the QAOA+ algorithm in Section II.3.
@@ II.1 Review of MIS
The MIS is defined on a graph $G=(V,E)$, where $V=\left\{{0,\dots,n-1}\right\}$ is the set of vertices and $E=\{\{u,v\}\mid u,v\in V,u\neq v\}$ is the set of edges. In the graph theory, two vertices $u$ and $v$ are considered adjacent if they are connected by an edge $(u,v)$, and a subset of vertices $V_{s}$ is called an independent set of $G$ if no two vertices in $V_{s}$ are adjacent. The MIS is the vertex set with the most number of vertices among all independent sets reducibility , and the number of vertices in the MIS is known as the independence number, denoted as $\beta(G)$. While the solution to the MIS problem for a graph may not be unique, the independence number is always uniform.
@@
Figure 2: A schematic of solving the MIS problem on the target graph using QAOA.
For any target graph $G$, the independent vertex subsets in an induced subgraph $G_{q}$ of $G$ are feasible solutions for $G$. That is, there is at least one induced subgraph $G_{q}=(V_{q},E_{q})$ that has an identical MIS solution to $G$, where $V_{q}\subseteq V$ and $E_{q}\subseteq E$. As illustrated in FIG. 1, the right graph is the induced subgraph of the left graph, and they have the same MIS solution $\left\{1,2,3\right\}$, but the MIS solution $\left\{1,2,4\right\}$ to the target graph is not the MIS solution of the right graph because the node $4$ is not in this induced subgraph, which implies that the MIS solution of the target graph may not always be a solution for its induced subgraph. In this paper, we denote the number of vertices and edges in $G$ as $n$ and $m$, respectively. To formalize the MIS problem, a binary variable $x_{u}$ for each vertex $u$ in $V$ is introduced, where $x_{u}=1$ represents that the vertex $u$ is in the vertex subset, otherwise, $x_{u}=0$. Each way of vertex partition corresponds to a unique bit string $x$, and there are $2^{n}$ forms for $x$ in total, where $x=x_{0}x_{1}\dots x_{n-1}$. The goal of the MIS problem is to maximize the following objective function
@@
when the independence constraint $\sum_{(u,v)\in E}x_{u}x_{v}=0$ is met. This constraint can be incorporated into the objective function using the Lagrange multiplier method Iterative_MIS , thus converting the constrained problem to an unconstrained problem. The obtained classical cost function is
@@
where the value of $\lambda$ is determined through additional hyper-parameter optimization steps QIRO , and $\lambda>1$ can impose penalties on those terms that violate independence constraints Ising_formulations , making the maximal cost function value correspond to the exact MIS solution instead of other feasible or infeasible bit strings. Here, $C_{\text{max}}$ equals the independence number $\beta(G)$. The MIS problem has been rigorously proven to be NP-hard. As the number of vertices in a graph increases, the time complexity of solving the MIS problem using exact algorithms, such as backtracking backtracking_algorithm_MIS and branch-and-bound methods branch_and_bound_MIS , grows exponentially exact_algorithm_MIS . This rapid escalation in computational demands renders exact algorithms increasingly impractical for large-scale instances. To address this challenge, researchers have proposed heuristic algorithms. Some of the representative classical heuristic approaches include evolutionary algorithms evolutionary_algorithm_MIS , tabu search tabu_search , and local search hybrid_iterated_LS ; iterated_LS ; local_search_MIS . Although these algorithms sacrifice some degree of precision compared with the exact algorithms, they may provide quasi-optimal solutions in a shorter time, striking a balance between efficiency and accuracy.
@@ II.2 Review of QAOA
QAOA is an approximate algorithm and provides an approximation solution to the solved problem. QAOA encodes the solution of the problem into the ground state of the target Hamiltonian $H_{C}$ and aims to approximate this ground state through a PQC qaoa . This evolution starts from the ground state $|s\rangle$ of the initial Hamiltonian $H_{B}$, with a quantum circuit built using $p$-level QAOA ansatz, where $p$ is the level depth. Notably, the initial state $|s\rangle$ is required to be trivial to implement. That is, $|s\rangle$ can be prepared by a constant-depth quantum circuit. For the original QAOA, one level QAOA ansatz consists of two variational QAOA parameters (denoted as $\gamma_{i}$ and $\beta_{i}$) and two unitaries $e^{-i\gamma_{i}H_{C}}$ and $e^{-i\beta_{i}H_{B}}$, where $i=1,2,\cdots,p$. Generally, the initial Hamiltonian is chosen as $H_{B}=-\sum_{u=0}^{n-1}\sigma_{u}^{x}$ because its ground state $|s\rangle=|+\rangle^{\otimes n}$ can be trivial to prepare. Here, $\sigma_{u}^{x}$ refers to the Pauli-X operator applied to the $u-$th qubit, effectively flipping its state. To encode the MIS solution into the ground state of $H_{C}$, the cost function $-C(x)$ is transformed into the target Hamiltonian
@@
by transforming each binary variable $x_{u}$ to a quantum spin $\frac{I-\sigma_{u}^{z}}{2}$ qaoa , where $\sigma_{u}^{z}$ refers that applying Pauli-Z to the $u$-th qubit. The ground state of $H_{C}$ can be approximately obtained by alternately applying unitaries operators $e^{-i\gamma_{i}H_{C}}$ and $e^{-i\beta_{i}H_{B}}$ to the initial quantum state $|s\rangle$. After applying these unitaries for $p$ levels, the output state of the PQC is
@@
where $\bm{\gamma_{p}}=(\gamma_{1},...,\gamma_{p})$ and $\bm{\beta_{p}}=(\beta_{1},...,\beta_{p})$ are $2p$ variational QAOA parameters. The expectation function value of $H_{C}$ in this variational quantum state is defined as
@@
which can be calculated by repeated measurements of the output quantum state using the computational basis. To search for the global optimal parameters
$(\bm{\gamma_{p}}^{\text{opt}},\bm{\beta_{p}}^{\text{opt}})=\arg\underset{(\bm{\gamma_{p}},\bm{\beta_{p}})}{\max}F(\bm{\gamma_{p}},\bm{\beta_{p}}),$ (6)
@@
multiple optimization runs tend to be required. In each run, the search is typically done by starting with some guesses of the initial parameters. Then, the quantum computer measures the output state to compute the expected value of the output state and delivers the relevant information to the classical computer. These parameters are optimized by the classical optimizer and then delivered to the quantum computer. The outer parameter optimization is repeated until it meets the termination condition, such as the maximal number of iterations or convergence tolerance $\varepsilon$ MLI . To quantify the quality of the QAOA solution (i.e., measure the difference between the output state of QAOA and the ground state of $H_{C}$), the approximation ratio (AR)
$r=\frac{F({\bm{\gamma_{p}}},{\bm{\beta_{p}}})}{F_{\text{max}}}$ (7)
@@
is introduced, where $F_{\text{max}}$ is the negative of the ground state energy of $H_{C}$, with equaling $C_{\text{max}}$. For small graph instances, $C_{\text{max}}$ can be obtained by brute force search. For large graphs, an approximation of $C_{\text{max}}$ can be obtained by executing multiple runs of the classical heuristic algorithm and selecting the largest independent set found across these runs. AR reflects how close the solution provided by QAOA is to the exact solution, $r\leq 1$, with the value of 1 as the exact solution. Finally, to help readers better understand the QAOA, its detailed process is shown in FIG. 2.